Suppose that Trevor is a resident of The Bad Place who enjoys hosting loud parties late at night. This annoys Trevor’s neighbor Chidi, who likes going to sleep early. Suppose that Trevor enjoys each additional night of partying at $100, while Chidi values a good night’s sleep at $200.

a) If there are no laws in The Bad Place limiting residents’ behavior, will Trevor host parties every night?

b) What is the socially efficient outcome in this situation? What transaction needs to take place for this outcome to be realised?

c) What happens if they need to pay a translator $150 in order to be able to communicate, as Chidi can only speak French?

Go back to the original case, and suppose that Trevor’s friends Vicki and Shawn also get enjoyment out of attending Shawn’s parties, while Chidi’s friends Eleanor and Janet would rather enjoy a quiet night in. Vicki and Shawn benefit from the parties $240 and $300 per night respectively. Eleanor and Janet enjoy the quiet $50 and $400 respectively.

d) What is the new socially optimal outcome? How will this outcome be achieved?

e) What does this tell us about the requirements for Coase Theorem to realise the efficient outcome?
Answer #1

a) Yes, As Trevor values each additional night of partying at $100, therefore he wll keep on hosting parties every night.

b) Socially efficient outcome will be where overall the society gets most value. If in case Trevor keeps hosting parties, then society gets $100 value. But if chidi pays Trevor $100 not to host parties, then total society value if $200 and that is the socially optimal outcome. The Transaction is chidi playing Trevor 100 not to host parties.

c) In any case. Trevor would need to get atleast $100 to stop partying. Since they can't communicate and need to pay 150 to a translator, the total cost of stopping a party will become 100+150= 250. Then chidi doesn't have that money because if she spends 250 to get a benefit of 200, its not worth of it. Hence the transaction wouldn't take place and the parties keep on continuing.

d) Total benefit with parties = 100+240+300 = 640.

Total benefit with sleep = 200+50+400 = 650.

-> As benefit with sleep is more than profits, chidi and friends will pay trevor and also friends not to party. That will be the socially optimal outcome as total social benefit in that case will be $650.

***AS PER HOMEWORKLIB POLICY, I HAVE ANSWERED FIRST FOUR QUESTIONS, PLEASE UPHOLD MY ANSWER.